# Auditing Human Decision-Making in High-Stakes Environments via Prescriptive AI: A Stress-Test on Real-Time Tactical Management

### v5

@@ 4.7.3 Formal Definition of Prescriptive AI
  • Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias BaronHershey1988.
In its lexical sense, a “prescription” is the action of laying down authoritative rules or directions, or more generally, something prescribed as a rule mw:prescription. By this definition, a prescriptive statement is inherently normative: it asserts what ought to be done rather than merely predicting or suggesting an outcome. Consequently, any system that cannot be inspected, audited, or contested cannot properly prescribe.
4.7.4 Decision-Making as Epistemic State Transition
The normative necessity of Prescriptive AI follows from a dynamic conception of rationality. In the framework of logical dynamics developed by van Benthem, decision-making is not primarily evaluated by realized outcomes, but by the epistemic state transitions induced by actions under partial and evolving information (vanBenthem2007; vanBenthem2011). Rationality, under this view, is a property of how an agent updates beliefs, constraints, and commitments at the moment of action, rather than of the stochastic realization of downstream consequences.
4.7.5 Prescriptive AI vs. Predictive and Outcome-Driven Systems
Dimension Predictive AI Prescriptive Analytics Prescriptive AI
Primary Objective Forecast events or actions Recommend actions to optimize objectives Support normative human decision-making
4.7.6 Boundary Cases Within Prescriptive Analytics
Optimization-based systems compute actions that maximize objective functions under constraints Bertsimas2020. While effective in structured settings, they typically assume stable objectives and do not audit human decision justification under uncertainty.
Reinforcement Learning systems prescribe actions by optimizing expected reward SuttonBarto2018. Despite their prescriptive nature, they are often opaque and automation-oriented, limiting contestability and interpretability, and are particularly vulnerable to outcome bias in stochastic environments.
4.8 Hybrid Systems and Overlapping Set Membership
4.8.1 Cross-Domain Applicability: From Pitch to Critical Operations
While elite-level soccer serves as an observable laboratory for adversarial decision-making, the structural constraints of the domain—namely irreversibility, severe time pressure, and asymmetric risk—are isomorphic to those encountered in several safety-critical sectors. The prescriptive auditing framework proposed herein passos2025 does not rely on domain-specific mechanics, but on the temporal auditing of decision-relevant performance signals and agent state evolution. Consequently, the architecture is transferable to environments in which human judgment is susceptible to fatigue, status quo bias, or desensitization under sustained cognitive load.
  • Aviation and Control Room Operations: In air traffic control and other safety-critical control rooms, operator fatigue and cognitive degradation constitute latent failure modes that are difficult to observe directly. A prescriptive auditor could assess indicators of cognitive performance degradation—such as response latency or interaction patterns—to recommend preemptive role rotation or task redistribution before safety margins are compromised. The “substitution” logic instantiated in this work thus generalizes to proactive management of human operational capacity.
Across these domains, the core function of the system remains invariant: to decouple the evaluation of the decision state at time $t$ from stochastic outcomes at time $t+n$, thereby mitigating outcome bias in high-variance, high-stakes environments.
4.8.2 Terminological Convergence and Alternative Taxonomies
This perspective aligns with foundational work on the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty (vanBenthem2007; vanBenthem2011). While the present work does not implement the formal machinery of dynamic epistemic logic, it adopts this conceptual stance by evaluating decisions based on the informational state and normative criteria available at the moment of action, rather than on the stochastic realization of downstream outcomes. By treating decision-making as a dynamic process of information uptake and evaluation, Prescriptive AI analytically decouples the agent’s reasoning from stochastic environmental variability, enabling principled auditing of decision quality independently of realized results.
4.8.3 Implications for High-Stakes Decision Making
By explicitly positioning Prescriptive AI as a normative subset of Prescriptive Analytics, and Prescriptive Analytics as a structured subset of a broader heterogeneous space of action-oriented systems, this work resolves a persistent source of conceptual confusion in both academic and industrial discourse Sun2020; Wissuchek2022.
This hierarchy clarifies that:
  • Not all action-oriented systems are prescriptive analytics;
  • Not all prescriptive analytics systems are Prescriptive AI;
  • Hybrid systems may legitimately belong to multiple conceptual sets.
4.9 Ethical Use of AI in High-Stakes Decision Systems
Auditing artificial intelligence systems in high-stakes decision-making contexts is not merely a technical challenge but a fundamentally ethical one. As argued by Rudin2019, deploying opaque, black-box models in domains where decisions have irreversible consequences is inherently problematic. When such systems are extended to environments with intense public scrutiny and continuous human judgment—such as elite professional soccer—the ethical risks are amplified rather than mitigated.
Epistemic opacity and public accountability.
Institutional distrust and risk of misuse.
Prescriptive AI as an ethical design principle.
This work adopts interpretability and prescriptiveness as core ethical design principles. By prioritizing transparency, auditability, and human agency, the proposed system aligns with emerging arguments that intrinsically interpretable models are not only preferable but necessary in domains where errors carry disproportionate consequences (rudin). Rather than automating authority, the system enhances human decision-making, offering a practical pathway for deploying AI responsibly in high-stakes environments.
@@ 5 Acknowledgments
In the spirit of Open Science, the complete source code, including pre-processing pipelines, fuzzy inference engines, and validation scripts, is publicly available at https://github.com/Pedro-Passos77/AI-Assisted-Substitution-Decisions-A-Fuzzy-Logic-Approach-to-Real-Time-Game-Management. This manuscript constitutes a substantially expanded version of the research accepted by the Wharton Sports Analytics Journal and the accepted in the AAAI 2026 Bridge on LMreasoning) Passos2025_Wharton passos2026, with improvements focused on methodological robustness, theoretical foundation, proposed paradigm and extended case study. We gratefully acknowledge Luca Pappalardo and collaborators for making the Soccer match event dataset publicly available (Pappalardo2019), which was essential for this study. Generative AI tools were used to assist in the drafting and structural organization of this manuscript under strict human supervision; the authors remain fully responsible for the final content and its normative justification.

### v6

@@ 7 Discussion
    Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias [4].
7.1 Behavioral Audit: Quantifying Cognitive Biases
The “Fagner Paradox” as Status Quo Bias.
The “Lukaku Paradox” as Salience Masking and Outcome Bias.
The validity of this auditing approach rests on a dynamic conception of rationality, grounded in the logical dynamics framework of van Benthem [48]. Under this view, rationality is a property of the epistemic update at the moment of commitment ($t$), rather than of the stochastic realization of outcomes at a later time ($t+n$). If decision quality is defined prior to outcome realization, then systems designed to support human judgment must operate strictly on the informational state available at the moment of action. This perspective justifies our rejection of outcome-based evaluation metrics: a decision is not “correct” because the team won, but because it constituted a normatively coherent response to the risk state at time $t$.
7.3 Implications for Human Decision-Making
  • 1.
  • 2.
7.4 Normative, Epistemic, and Institutional Implications of Prescriptive AI
7.4.1 Normative Scope and Limitations
A prescriptive audit is only as meaningful as the normative assumptions it encodes. The framework proposed here deliberately commits to an explicit model of risk, temporal aggregation, and asymmetric costs, enabling transparent evaluation of decision quality at the moment of commitment. This explicitness, however, also implies that prescriptive recommendations are inherently contingent rather than universally binding. Alternative normative models—reflecting different institutional priorities, ethical trade-offs, or risk tolerances—would yield different but equally auditable prescriptions.
Crucially, this contingency does not weaken the prescriptive paradigm; rather, it constitutes its primary epistemic advantage. By making normative assumptions explicit and contestable, Prescriptive AI avoids the false objectivity often implied by outcome-optimized or imitation-based systems. Disagreement over normative choices is therefore not a failure of the system, but an expected and desirable feature of a transparent decision audit.
7.4.2 Prescriptive Auditing versus Predictive Alignment
It is important to distinguish prescriptive auditing from predictive alignment objectives commonly pursued in contemporary AI systems. Alignment-based approaches seek to minimize divergence between machine outputs and observed human behavior, implicitly treating historical expert decisions as normative ground truth. In contrast, prescriptive auditing explicitly allows—and indeed foregrounds—systematic disagreement between human judgment and normative risk assessment.
The objective of Prescriptive AI is not convergence between human and machine, but diagnosability of decision-making processes. Persistent divergence identifies regions in which human heuristics, institutional pressures, or cognitive biases override normatively justified responses. From this perspective, disagreement is not an error signal to be minimized, but an epistemic signal to be analyzed.
7.4.3 Robustness under Distributional Shift
A central limitation of predictive decision-support systems is their sensitivity to distributional shift. Models trained to forecast outcomes or imitate historical behavior implicitly assume stability in both the environment and expert competence. When these assumptions fail—as they routinely do in high-variance, adversarial domains—predictive performance degrades and interpretability diminishes.
Prescriptive auditing offers a fundamentally different robustness profile. Because recommendations are derived from explicit normative criteria applied to the current epistemic state, their validity does not depend on the stationarity of empirical distributions. While statistical estimates may fluctuate, the criteria governing decision evaluation remain stable. This property is particularly important in dynamic institutional settings, where tactics, incentives, and environmental conditions evolve faster than predictive models can be reliably retrained.
7.4.4 Human–AI Disagreement as an Epistemic Signal
The systematic disagreements observed between the prescriptive system and human decision-makers constitute a primary object of scientific interest. Rather than treating these divergences as isolated errors, the prescriptive framework enables their localization, quantification, and temporal analysis. Disagreement patterns reveal where human agents systematically discount cumulative risk, overweight salient events, or defer action due to heuristic inertia.
This reframing elevates human–AI interaction from an optimization problem to an epistemic diagnostic tool. The goal is not to replace human judgment, but to expose its structural limitations under uncertainty. In this sense, prescriptive systems function less as decision-makers and more as instruments for auditing the rational coherence of human action.
7.4.5 Prescriptive Auditing versus Causal Attribution
Prescriptive auditing must be clearly distinguished from causal inference and counterfactual explanation. The framework does not seek to explain why outcomes occurred, nor to assign causal responsibility to individual decisions. Instead, it evaluates whether a decision was normatively justified given the information available at the moment of commitment.
By operating strictly on epistemic states rather than outcome-generating mechanisms, prescriptive systems avoid retrospective rationalization. A decision is not deemed incorrect because it led to an unfavorable outcome, nor correct because it coincided with success. This separation is essential for maintaining normative coherence in environments dominated by stochastic variance.
7.4.6 Institutional Scalability and Governance
While the present analysis focuses on individual expert decisions, the prescriptive paradigm naturally extends to institutional contexts. Aggregated audit trails enable organizations to identify recurrent blind spots, structural incentives for inaction, and systematic deviations from stated risk policies. Unlike outcome-based evaluations, which conflate skill with luck, prescriptive records permit longitudinal assessment of decision quality independent of stochastic realization.
This property has direct implications for governance in domains subject to public accountability. In settings where decisions must be justified to regulators, stakeholders, or courts, prescriptive auditing provides a defensible record of rational deliberation grounded in the information available at the time of action.
7.5 Ethical Accountability and Epistemic Trust
Resolving Epistemic Opacity.
The Audit Trail as Institutional Defense.
7.6 The Prescriptive Paradigm and the Future of AI Alignment
Current debates on AI alignment largely focus on aligning model outputs with human preferences or values, typically through techniques such as reinforcement learning from human feedback (RLHF) or constitutional AI [1, 3]. These approaches implicitly assume that historical human judgments constitute valid ground truth for training objectives.
The Prescriptive AI paradigm challenges this assumption fundamentally. Rather than treating human decisions as labels to be replicated, prescriptive systems treat them as objects of audit. This inversion has profound implications for AI safety and governance:
Alignment as Auditing, Not Imitation.
In domains where human experts exhibit systematic biases—as demonstrated empirically in this work—alignment through imitation perpetuates rather than corrects decision errors. The Imitation Incompleteness Theorem formalizes this limitation: without external normative signals, supervised learning cannot escape the structural biases embedded in training data.
Prescriptive AI reframes alignment as a bidirectional process: rather than aligning AI to human behavior, the system surfaces divergences that allow humans to align their decisions to normative criteria. This constitutes a form of normative bootstrapping, in which human and machine co-evolve toward improved decision quality.
Implications for High-Stakes AI Deployment.
As AI systems are deployed in safety-critical contexts—autonomous vehicles, medical diagnosis, financial regulation—the tension between imitation and prescription becomes acute. A self-driving car trained solely on human driving behavior will reproduce human errors (e.g., delayed reaction to hazards). A prescriptive system, by contrast, evaluates driving decisions against explicit safety criteria, independent of typical human response times.
This distinction suggests that AI alignment objectives must be domain-dependent: in creative or preference-driven tasks, alignment through imitation may be appropriate; in high-stakes, normatively constrained domains, prescriptive auditing is essential.
The Prescriptive-Predictive Frontier.
Future research must delineate the boundary conditions under which each paradigm applies. We conjecture that prescriptive systems are necessary when:
  1. 1.
    Decisions are irreversible with asymmetric consequences,
  2. 2.
    Historical decisions encode systematic biases,
  3. 3.
    Normative criteria can be formalized independently of outcomes,
  4. 4.
    Human accountability requires explicit justification.
Conversely, predictive systems remain appropriate when human preferences constitute valid optimization targets and environmental dynamics are sufficiently stable.
Establishing this frontier formally—potentially through a generalization of Imitation Incompleteness Theorem to multi-agent or multi- objective settings—constitutes a critical research agenda for trustworthy AI.
7.7 Prescriptive AI and Human Expertise: Augmentation, Not Replacement
A recurring concern in human-AI collaboration research is the risk of deskilling: as systems assume cognitive tasks, human experts lose proficiency, creating long-term dependency [7, 2]. Prescriptive AI presents a distinct profile in this regard.
Preserving Deliberative Capacity.
Unlike fully automated systems, prescriptive auditing retains human agency at the point of commitment (Axiom 4: Contestability). The coach can—and frequently does—override algorithmic recommendations. This preserves the exercise of judgment even when augmented by computational support.
Empirical evidence from our case studies supports this: in 3 of 4 substitutions (Willian, Jesus, Paulinho), the human decision aligned with the system. In 1 case (Fagner), the human overrode the system—demonstrating that decision authority remained with the expert. This is fundamentally different from automation, where human input is reduced to supervisory monitoring.
Skill Transformation, Not Atrophy.
Rather than replacing expertise, prescriptive systems transform the skill profile required:
  • 1.
    From pattern recognition to norm evaluation: Experts shift from detecting performance decay (now automated) to evaluating whether system recommendations align with strategic context.
  • 2.
    From reactive to proactive: Alerts surface latent risks before they become salient, enabling anticipatory rather than reactive intervention.
  • 3.
    From intuition to justification: Decisions must be explained (to stakeholders, media, institutions), elevating the role of explicit reasoning over tacit intuition.
This skill transformation is analogous to how calculators shifted mathematical expertise from computation to problem formulation. The cognitive labor remains, but its locus changes.
Risk of Over-Reliance: Automation Bias Redux.
Despite preserving agency, prescriptive systems remain vulnerable to automation bias [32]: humans may defer excessively to algorithmic recommendations, even when contextual factors warrant override.
Our framework mitigates this through:
  1. 1.
    Transparency: Rules and thresholds are inspectable,
  2. 2.
    Contestability: Overrides are explicitly documented,
  3. 3.
    Audit trails: Post-hoc review of decisions vs. recommendations.
Longitudinal field studies are needed to assess whether these safeguards suffice in practice. We hypothesize that over-reliance risk correlates with:
  • 1.
    System accuracy (higher accuracy $\rightarrow$ stronger deference),
  • 2.
    Organizational culture (hierarchical $\rightarrow$ less contestation),
  • 3.
    Temporal pressure (urgency $\rightarrow$ default to algorithm).
The Role of Training.
Effective deployment of prescriptive AI requires training decision-makers in system critique, not system operation. Experts must learn to:
  • 1.
    Identify edge cases where rules may misfire,
  • 2.
    Recognize when contextual factors (invisible to the system) override statistical signals,
  • 3.
    Calibrate confidence in recommendations based on situational uncertainty.
This represents a new pedagogical challenge: teaching humans to collaborate with AI as critical auditors, not passive consumers. Educational programs in medicine, finance, and other high-stakes domains must integrate training in prescriptive system literacy.
@@ 8 Acknowledgments
In the spirit of Open Science, the complete source code—including all pre-processing pipelines, fuzzy inference engines, and validation scripts—is publicly available at https://github.com/Pedro-Passos77/AI-Assisted-Substitution-Decisions-A-Fuzzy-Logic-Approach-to-Real-Time-Game-Management . This manuscript constitutes a substantially expanded and revised version of prior research accepted by the Wharton Sports Analytics Journal and by the AAAI 2026 Bridge on LM Reasoning [34, 35]. The present work introduces significant extensions in methodological robustness, theoretical grounding, the proposed decision-making paradigm, and an expanded case study.
We gratefully acknowledge Luca Pappalardo and collaborators for making the Soccer Match Event Dataset publicly available [30], which was essential to this study. The authors declare no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper, neither this research has received any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Generative AI tools (Gemini 3 Pro and ChatGPT 5.2) were used to assist with programming tasks and the structural organization of the manuscript under strict human supervision. The authors retain full responsibility for the final content, interpretations, and normative justification of the proposed framework.